# Curriculum Negative Mining For Temporal Networks

### v1

@@ 4.1 Curriculum Learning
4.2 Negative Pool
The negative pool, which is an essential component of the curriculum learning strategy, is also constructed dynamically.
To ensure efficiency, we start the entire process by sampling a negative pool $V^{-}_{u}$ of fixed size $M$ for each positive sample $(u,v^{+},t)$. During the initial epochs, we sample $V^{-}_{u}$ as a mixture of historical neighbors and random negatives, a simple approach to prevent overcrowding with easy negatives. Random samples here serve a different purpose than before and are included to capture newly emerging preferences not disclosed by users’ historical interactions. This process can be written as
$V^{-}_{u}=V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (4)
Once the model learns the fundamentals, we raise the difficulty level. We introduce this advanced stage when sampling proportion $\pi_{e}$ falls below a preset threshold $\tau$ for the first time. Since $\pi_{e}$ decreases only if the model is learning well, this condition ensures that the model only encounters more challenging negatives after it has adequately learned the basics. The indicator $\mathcal{C}_{\pi_{e}}$ is determined as follows:
$\mathcal{C}_{\pi_{e}}=\mathbbm{1}(\pi_{e}\leq\tau).$ (5)
In this advanced stage, we implement a hard negative cache $V^{\text{hard}}_{u}$ of size $\frac{M}{2}$ to store a subset of $V^{-}_{u}$ from the current epoch, focusing on likely false positives and those showing less learning in previous epochs. The cached negatives are sampled proportionally according to the negatives’ performance metrics, as outlined in Eq. 6.
$V^{\text{hard}}_{u}\sim\{p_{uv}-\alpha_{e}\times\text{std}(P_{uv})|\forall v\in V^{-}_{u}\},$ (6)
where $|V^{\text{hard}}_{u}|=\frac{M}{2}$, $p_{uv}$ denotes the probability score of node pair $(u,v)$ at the current epoch, $P_{uv}$ records the historical scores within the last five epochs, and std($\cdot$) is set to 00 when results are not available. $\alpha_{e}$ is a hyper-parameter that controls the importance of prediction standard deviation at the $e$-th epoch. To manage potential instability in early learning stages, we initially assign a modest weight to the standard deviation and gradually increase it. Specifically, we use $\alpha_{e}=\alpha_{\text{max}}\times\min(\frac{e}{E^{\prime}},1)$, where $E^{\prime}$ is the epoch after which the weight remains constant, and $\alpha_{\text{max}}$ is the maximum weight that standard deviations could contribute to the performance metrics.
This hard negative cache then accounts for half of the negative pool in the next epoch, while the remainder continues to be a mix of historical neighbors and random samples, maintaining the previous proportion. Therefore, the new updating strategy for each negative pool $V^{-}_{u}$ is written as
$V^{-}_{u}=V^{\text{hard}}_{u}\cup V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (7)
4.3 Negative Selection
The second critical component of our framework is the selection function used to choose negative samples from the negative pool $V^{-}_{u}$. This process consists of two steps: disentangle irrelevant factors of negative nodes and score negative samples with time embeddings.
Irrelevant Factor Disentanglement
To facilitate TGNNs to differentiate positive from negative samples, we disentangle nodes’ relevant and irrelevant factors, which is in line with DENS [17].
For each positive sample $(u,v^{+},t)$, we construct a positive gate. This gate leverages the embeddings of the source node $h_{u}$ and the destination node $h_{v^{+}}$ to isolate the relevant factor $h_{v^{+}}^{r}$ of the destination node $v^{+}$, illustrated as
$h_{v^{+}}^{r}=h_{v^{+}}\odot\sigma(W_{\text{src}}h_{u}+W_{\text{dst}}h_{v^{+}}+b_{p}),$ (8)
where $W_{\text{src}}$, $W_{\text{dst}}$ and $b_{p}$ are trainable parameters shared by all positive samples, $\sigma$ is the sigmoid function, and $\odot$ is the element-wise product. This factor and the embedding $h_{v^{-}}$ of the corresponding candidate negative $v^{-}\in V^{-}_{u}$ are then used as inputs to isolate the relevant factor $h_{v^{-}}^{r}$ of the negative.
$h_{v^{-}}^{r}=h_{v^{-}}\odot\sigma(W_{p}h_{v^{+}}^{r}+W_{n}h_{v^{-}}+b_{n}),$ (9)
where $W_{p}$, $W_{n}$, and $b_{n}$ are trainable parameters shared by all candidate negatives in the negative pool. Factors other than relevant factors are denoted as irrelevant factors:
$h_{v^{+}}^{ir}=h_{v^{+}}\ominus h_{v^{+}}^{r},\quad h_{v^{-}}^{ir}=h_{v^{-}}\ominus h_{v^{-}}^{r},$ (10)
where $h_{v^{+}}^{ir}$ and $h_{v^{-}}^{ir}$ denote irrelevant factors of the destination node $v^{+}$ and the candidate negative $v^{-}$, respectively, and $\ominus$ is element-wise subtraction.
where $s_{g1,g2}^{sgn1,sgn2}=MLP(h_{u},h_{v^{sgn1}}^{g1})-MLP(h_{u},h_{v^{sgn2}}^{g2})$ is the scoring function between different factors, with $MLP(\cdot,\cdot)$ computing the probability score from the given embeddings, and a minus sign is added to perform maximization.
Temporal Scoring Function
As our goal is for the model to predict neighbors based on relevant factors, the most useful negatives for training are those that differ from the positives only in their relevant factors. In practice, this involves selecting negative samples from the negative pool $V^{-}_{u}$ by maximizing the differences in relevant factors and minimizing the differences in irrelevant factors. This approach is written as
$s_{uv^{-}}$ $=$ $\beta_{e}\times s_{r,r}^{+,-}-(1-\beta_{e})\times s_{ir,ir}^{+,-},$ (12)
where $s_{uv^{-}}$ denotes the scoring function and $\beta_{e}$ is a hyper-parameter that balances the contributions of relevant and irrelevant factors at epoch $e$. Intuitively, we want the model to first learn which factors are important and then understand how these factors drive interactions. This entails initially focusing on aligning irrelevant factors and later shifting to differentiating relevant factors. Specifically, we set $\beta_{e}=\min(\frac{e}{E^{\prime\prime}},1)$, where $E^{\prime\prime}$ marks the epoch from which the score is solely determined by relevant factors.
Additionally, to capture the unique aspects of temporal networks, we incorporate three Time2Vec models to represent complex temporal patterns, including recurrent and periodic behaviors. Therefore, the final temporal scoring function used for selecting negative samples is written as
$\hat{s}_{uv^{-}}=t_{u,v^{+}}\ast(t_{u,v^{-}}\odot t_{v^{-}})\odot s_{uv^{-}},$ (13)
where $t_{u,v^{+}}$, $t_{u,v^{-}}$, and $t_{v^{-}}$ are Time2Vec embeddings for the timestamps of the positive sample, the interaction between $u$ and the candidate negative $v^{-}\in V^{-}_{u}$, and the negative’s most recent occurrence, respectively. Default inputs of 0 are provided into Time2Vec for cases where the negative is not a historical neighbor or if it never occurs. By incorporating the product of these three embeddings, we seek to prioritize historical neighbors and recently popular nodes. The rationale is that these negatives possess attributes that were favored by the source nodes not long ago, and they might trick the model. So, by assigning more weight to these negatives – particularly those that differ from positives only in relevant factors – we can enhance the precision of the model’s decision boundary and reduce the occurrence of false negatives.
@@ 5.1 Datasets and Evaluation
Table 1: Statistical Overview of the Datasets. Average Proximity is the average step difference between nodes and their historical neighbors, Recency is the percentage of current interactions that replicate the last, Repetition is the proportion of repeated edges, $d_{\text{max}}$ represents the maximum temporal node degree.
@@
MOOC Interaction No 7047 411749 30 20% 57% 4.80E-08 1.95E+04 1 month
@@
We use average precision (AP) as our primary evaluation metric. Specifically, we compare AP for each dataset across all combinations of TGNNs and sampling methods, since typically only the optimal setup is used in practice. Additionally, given that we conduct experiments across 12 datasets, we evaluate the overall performance of each sampling method using average rank.
@@ 5.2 Experiment Setting
Baselines
We compare our method against two commonly adopted negative sampling strategies – random sampling and recent sampling [24, 33] – along with ENS [7], the only existing method known to us tailored for temporal networks. Random sampling selects negative samples based on a uniform distribution. Recent sampling selects the latest historical neighbors as negatives, offering a straightforward approach to addressing issues of positive sparsity and positive shift. Finally, ENS trains models using a predetermined proportion of random negatives and historically appearing nodes while ensuring that the average temporal proximity reaches a specified level.
Implementation Details
@@
The contrastive loss coefficient $\lambda$ and the $L_{2}$ regularization coefficient $\mu$ are set to ensure a balanced treatment across different loss components. The remaining hyperparameters are carefully tuned through grid search. The size of the negative pool $M$ is tested within $\{4,8,12,16\}$; the proportion of historical neighbors $\gamma_{\text{hist}}$ within $\{0.3,0.4,0.5,0.6,0.7\}$; and the hard negative cache’s initiation threshold $\tau$ within $\{0.3,0.5,0.7\}$. The stopping epoch $E^{\prime}$ and the maximum weight $\alpha_{\text{max}}$ for $\alpha_{e}$ are explored over $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0,0.006,0.012,0.024\}$ respectively, where $E_{0}$ denotes the maximum number of epochs. Similarly, $E^{\prime\prime}$ for $\beta_{e}$ and $\delta_{\text{min}}$ for $\delta_{e}$ are searched within $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0.3,0.5,0.7\}$. All chosen values are included in the code. We will elaborate on tests involving $\alpha_{e}$, $\beta_{e}$, $\pi_{e}$, and $\delta_{e}$ in a later section.
@@ 5.3 Experimental Results
As shown in Table 2, our method consistently achieves the best ranks in both historical and inductive settings, underscoring its efficacy.
@@
Table 2: Average Precision of link prediction across 12 datasets and three TGNNs. bold denotes the best performance achieved for each dataset, - indicates failed results, Avg Rank refers to the average ranks of each sampling method.
Dataset Model Transductive Inductive
Random Recent ENS CurNM Random Recent ENS CurNM
uci DySAT 0.763 0.787 0.763 0.764 0.726 0.727 0.737 0.741
JODIE 0.839 0.564 0.855 0.848 0.636 0.542 0.648 0.640
TGN 0.824 0.651 0.849 0.832 0.677 0.662 0.700 0.695
USLegis DySAT 0.505 0.602 0.564 0.539 0.480 0.584 0.510 0.517
JODIE 0.492 0.527 0.501 0.515 0.597 0.596 0.618 0.620
TGN 0.808 0.797 0.833 0.868 0.763 0.730 0.789 0.809
CanParl DySAT 0.586 0.539 0.597 0.577 0.672 0.630 0.674 0.664
JODIE 0.572 0.569 0.570 0.652 0.533 0.539 0.530 0.570
TGN 0.747 0.705 0.713 0.791 0.626 0.624 0.604 0.650
enron DySAT 0.744 0.727 0.742 0.742 0.704 0.698 0.703 0.706
JODIE 0.800 0.752 0.801 0.826 0.724 0.693 0.716 0.739
TGN 0.771 0.842 0.824 0.848 0.730 0.811 0.769 0.787
WIKI DySAT 0.836 0.844 0.843 0.887 0.821 0.804 0.823 0.834
JODIE 0.791 0.734 0.786 0.818 0.701 0.701 0.712 0.728
TGN 0.906 0.883 0.908 0.917 0.856 0.837 0.860 0.868
MOOC DySAT 0.699 0.611 0.720 0.710 0.684 0.592 0.686 0.681
JODIE 0.670 0.564 0.679 0.679 0.580 0.574 0.576 0.577
TGN 0.764 0.621 0.774 0.767 0.688 0.653 0.663 0.690
UNtrade DySAT 0.527 0.534 0.528 0.527 0.538 0.556 0.539 0.542
JODIE 0.609 0.655 0.603 0.613 0.609 0.680 0.610 0.619
TGN 0.644 0.660 0.647 0.665 0.661 0.665 0.660 0.681
REDDIT DySAT 0.835 0.813 0.842 0.840 0.823 0.814 0.829 0.819
JODIE 0.823 0.781 0.835 0.829 0.787 0.777 0.794 0.797
TGN 0.878 0.848 0.890 0.878 0.874 0.853 0.871 0.877
myket DySAT 0.668 0.738 0.665 0.667 0.596 0.689 0.592 0.596
JODIE 0.674 0.664 0.685 0.688 0.530 0.533 0.545 0.537
TGN 0.694 0.679 0.692 0.702 0.582 0.539 0.581 0.588
UNvote DySAT 0.510 0.502 0.508 0.516 0.511 0.499 0.509 0.519
JODIE 0.684 0.639 0.696 0.714 0.694 0.635 0.698 0.703
TGN 0.681 0.601 0.693 0.709 0.650 0.594 0.660 0.673
LASTFM DySAT 0.701 0.640 0.704 0.754 0.709 0.622 0.712 0.738
JODIE 0.536 0.549 0.551 0.558 0.549 0.526 0.558 0.571
TGN 0.742 0.603 0.740 0.700 0.706 0.596 0.701 0.676
Flights DySAT 0.737 0.699 - 0.511 0.710 0.689 - 0.521
JODIE 0.736 0.484 - 0.772 0.656 0.605 - 0.689
TGN 0.771 0.573 - 0.807 0.696 0.665 - 0.747
Avg Rank 2.75 3.19 2.18 1.72 2.64 3.28 2.39 1.56
Transductive Setting
In the transductive setting, our method delivers the best results across 8 datasets, with its overall average rank nearly half a point better than the second-best method, ENS, and more than a point ahead of random and recent samplings. This suggests that our method outperforms all existing methods in predicting historical recurrence while maintaining strong performance in predicting random samples.
At the dataset level, our method excels in relatively dense networks with frequent recurrence of recent historical interactions (low Average Time Proximity), such as USLegis, CanParl, and Flights, where it surpasses the best existing methods by around 5%. This demonstrates that our method, by utilizing temporal-aware disentangled factors for selection, can more accurately distinguish between enduring and transient relationships over time. In networks with high repetition and recency effects – typically favorable for recent sampling – such as enron, UNtrade, and UNvote, our method still outperforms recent sampling. Furthermore, it achieves an average improvement of nearly 5% over the best random sampling and nearly 3% over ENS, showcasing its effectiveness in capturing periodic and recurrent patterns. These results confirm our method’s capability in addressing positive shift.
In very sparse networks, our method also outperforms random sampling across all datasets and, on average, surpasses ENS and recent sampling. Even where it does not lead, it remains competitively close to the top performer. The only exception is myket, whose sparsity, recency, and repetition make it akin to a static network, rendering ENS and CurNM – methods designed for temporal networks – less effective. Despite this, the overall performance indicates that our method successfully circumvents the steep gradient curve prevalent in hard negative sampling, effectively addressing positive sparsity even in less favorable scenarios.
Inductive Setting

### v2

@@ 5.1 Datasets
Table 1: Statistical Overview of the Datasets. “Average Proximity” is the average step difference between nodes and their historical neighbors, “Recency” is the percentage of current interactions that replicate the last, “Repetition” is the proportion of repeated edges, “$d_{\text{max}}$” represents the maximum temporal node degree.
@@
MOOC Interaction Yes 7047 411749 30 20% 57% 4.80E-08 1.95E+04 1 month
@@
5.2 Baselines and Evaluation Metrics
Baselines
We compare our method against two commonly adopted negative sampling strategies: random sampling and recent sampling [30, 39]. We also compare with ENS [10], the only existing method known to us that is specifically designed for temporal networks, and TASER [7], which focuses on positive sample selection.
  • 1.
    Random sampling selects negative samples based on a uniform distribution.
  • 2.
    Recent sampling selects the latest historical neighbors as negatives, offering a straightforward approach to addressing issues of positive sparsity and positive shift.
  • 3.
    ENS samples a fixed proportion of random negatives and historical neighbors while ensuring that the average temporal proximity of the selected negatives meets a predefined threshold.
  • 4.
    TASER, though primarily designed for denoising, includes a positive sample selection module. In this module, it selects positive samples with the highest predicted scores to reduce noise during training, while adding a small amount of noise to retain potentially informative but less confident positives. In our comparison, we adopt only this component.
@@ 5.2 Baselines and Evaluation Metrics
We use Average Precision (AP) as our primary evaluation metric. Specifically, we compare the AP of CurNM with the average AP of all baseline methods for each dataset-TGNN pair. We also compare all combinations of sampling methods and TGNNs for each dataset, since in practice, typically only the optimal setup is used. Additionally, given that our experiments span 12 datasets, we evaluate the overall performance of each sampling method using average rank.
@@ 5.3 Experiment Setting
The contrastive loss coefficient $\lambda$ and the $L_{2}$ regularization coefficient $\mu$ are set to ensure balanced contributions across different loss components and are used without additional tuning. The remaining hyperparameters are carefully tuned through grid search. The size of the negative pool $M$ is tested within $\{4,8,12,16\}$; the proportion of historical neighbors $\gamma_{\text{hist}}$ within $\{0.3,0.4,0.5,0.6,0.7\}$; and the hard negative cache’s initiation threshold $\tau$ within $\{0.3,0.5,0.7\}$. The stopping epoch $E^{\prime}$ and the maximum weight $\alpha_{\text{max}}$ for $\alpha_{e}$ are explored over $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0,0.006,0.012,0.024\}$ respectively, where $E_{0}$ denotes the maximum number of epochs. Similarly, $E^{\prime\prime}$ for $\beta_{e}$ and $\delta_{\text{min}}$ for $\delta_{e}$ are searched within $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0.3,0.5,0.7\}$. All chosen values are included in the code.
@@ 5.4 Experimental Results
As shown in Tables 2 and 4, our method consistently achieves the best ranks in both historical and inductive settings, underscoring its efficacy.
@@
Table 2: Average Precision of link prediction across 12 datasets and three TGNNs in the transductive setting. “AvgImp” denotes the average percentage improvement of CurNS compared to other negative sampling methods. Cells highlighted in red indicate that CurNS achieved a statistically significant improvement, while blue cells indicate a significant deterioration. Uncolored cells reflect no significant difference. NaN indicates failed results. “Avg Rank” refers to the average ranks of each sampling method.
Dataset Model Random Recent ENS TASER CurNM AvgImp (%)
uci DySAT 0.763 0.787 0.763 0.759 0.767 -0.140
JODIE 0.839 0.564 0.855 0.792 0.888 19.773
TGN 0.824 0.651 0.849 0.772 0.847 10.517
USLegis DySAT 0.505 0.602 0.564 0.566 0.557 0.071
JODIE 0.492 0.527 0.501 0.657 0.494 -8.007
TGN 0.808 0.797 0.833 0.699 0.870 11.412
CanParl DySAT 0.586 0.539 0.597 0.612 0.578 -0.735
JODIE 0.572 0.569 0.570 0.793 0.627 2.310
TGN 0.747 0.705 0.713 0.751 0.817 12.204
enron DySAT 0.744 0.727 0.742 0.731 0.734 -0.259
JODIE 0.800 0.752 0.801 0.553 0.788 11.006
TGN 0.771 0.842 0.824 0.763 0.877 9.855
WIKI DySAT 0.836 0.844 0.843 0.836 0.884 5.229
JODIE 0.791 0.734 0.786 0.773 0.803 4.245
TGN 0.906 0.883 0.908 0.888 0.921 2.742
MOOC DySAT 0.699 0.611 0.719 0.711 0.670 -1.767
JODIE 0.670 0.564 0.677 0.507 0.657 10.242
TGN 0.764 0.621 0.772 0.704 0.782 10.208
UNtrade DySAT 0.527 0.534 0.528 0.509 0.530 1.060
JODIE 0.609 0.655 0.611 0.600 0.616 -0.360
TGN 0.644 0.660 0.650 0.596 0.662 4.009
REDDIT DySAT 0.835 0.813 0.842 0.838 0.836 0.467
JODIE 0.823 0.781 0.834 0.839 0.835 1.960
TGN 0.878 0.848 0.890 0.881 0.881 0.818
myket DySAT 0.668 0.738 0.665 0.671 0.669 -2.171
JODIE 0.674 0.664 0.683 0.638 0.683 2.756
TGN 0.694 0.679 0.696 0.689 0.698 1.256
UNvote DySAT 0.510 0.502 0.508 0.513 0.516 1.486
JODIE 0.684 0.639 0.696 0.530 0.695 10.307
TGN 0.681 0.601 0.693 0.533 0.700 12.901
LASTFM DySAT 0.701 0.640 0.702 0.677 0.763 12.290
JODIE 0.536 0.549 0.552 0.537 0.570 4.862
TGN 0.742 0.603 0.740 0.583 0.716 8.778
Flights DySAT 0.737 0.699 nan 0.728 0.501 -30.482
JODIE 0.736 0.484 nan 0.715 0.782 25.814
TGN 0.771 0.573 nan 0.742 0.815 19.236
Avg Rank 3.11 3.75 2.36 3.53 2.03
5.4.1 Transductive Setting
In the transductive setting, our method consistently outperforms existing approaches across most dataset-model pairs, achieving an overall average rank that is 0.33 points better than the second-best method, ENS, and more than one full point ahead of random, recent, and TASER negative sampling strategies. This indicates that our method excels not only at capturing historical recurrence but also maintains strong generalization on randomly sampled negatives.
At the model level
CurNM consistently outperforms the mean of existing methods across all 12 datasets when paired with TGN. On average, it achieves a performance improvement of 8.7%.
For JODIE, CurNM also performs well overall, achieving an average improvement of 7.1%. The only statistically significant decline is observed on USLegis. In this case, since JODIE already struggles even before CurNM, reducing training difficulty - rather than increasing it - may be more effective, as evidenced by TASER’s strong performance.
CurNM shows comparatively weaker performance with DySAT. As a discrete-time TGNN, DySAT relies on static, snapshot-based modeling and lacks the temporal resolution needed for CurNM to operate effectively. This limitation is particularly evident in datasets such as myket, MOOC, and CanParl, where interactions are rarely repeated. Without an explicit memory mechanism, DySAT struggles to capture meaningful historical patterns, particularly those from further back in time, leaving CurNM with fewer temporal signals to leverage. The case of Flights further illustrates this issue: while CurNM selects hard negatives based on temporal similarity, DySAT’s snapshot aggregation collapses temporal order. Combined with the highly structured and repetitive nature of flight schedules, this increases the likelihood that selected negatives are actually temporally adjacent positives, raising the risk of false negatives. Nevertheless, our method still outperforms random sampling in most DySAT settings, demonstrating its robustness even under less favorable conditions.
Table 3: Percentage improvement of the best-performing CurNM method over the best-performing baseline (across all method–TGNN combinations) for each dataset.
Dataset Historical (%) Inductive (%)
uci 3.86 1.49
USLegis 4.44 3.55
CanParl 3.03 -1.19
enron 4.16 1.97
WIKI 1.43 1.74
MOOC 1.30 -1.69
UNtrade 0.30 0.44
REDDIT -1.01 0.00
myket -5.42 -13.35
UNvote 0.57 -1.29
LASTFM 2.83 5.50
Flights 5.71 5.77
At the dataset level
In Table 3, we compare the best-performing CurNM-model combination with the strongest baseline across all other method-model combinations to assess its practical effectiveness on each dataset. This reflects a realistic use case in which practitioners would pair CurNM with the most suitable model to achieve optimal performance.
CurNM performs particularly well on datasets characterized by frequent recurrence of historical interactions and relatively low recency, such as uci, USLegis, enron, and Flights, surpassing the best existing methods by approximately 3-6%. This proves that our method, through the incorporation of Time2Vec-based temporal modeling, can capture complex temporal dependencies, including long-term patterns that recent sampling fails to detect. Meanwhile, in networks with high recency - typically favoring recent sampling - such as WIKI, UNtrade, and UNvote, CurNM still outperforms the strongest baselines, further confirming its ability to capture positive shift.
CurNM also shows solid performance on user behavior-driven datasets with non-uniform and evolving temporal patterns, such as CanParl and LASTFM, achieving improvements of nearly 3%. This is expected, as our temporally-aware disentanglement mechanism is designed to capture individual-level preference shifts over time.
Finally, in highly sparse networks like MOOC and REDDIT, CurNM also achieves performance comparable to the best baselines. The only exception is myket, where recent sampling performs particularly well with DySAT. This is because users in myket often interact with multiple apps in short bursts but rarely return to the same app, leading to low repetition but strong temporal locality. As DySAT operates on discrete snapshots without memory, recent sampling aligns better with its structure by capturing relevant negatives and avoiding the noise introduced by historical negatives, which are less meaningful in this cold-start-heavy setting. Nevertheless, aside from this specific case, CurNM consistently matches all other methods. This demonstrates that our approach effectively circumvents the steep gradient issues common in hard negative sampling and successfully addresses positive sparsity.
5.4.2 Inductive Setting